# Absolute continuity of non-Gaussian and Gaussian multiplicative chaos measures

### v1

@@ 3.2.1. Reduction, and overview of the proof
in such a way that any $w_{1}\neq w_{2}$ in $\mathcal{N}_{n,j}(v)$ satisfies $1/8\geq|t_{n}(w_{1})-t_{n}(w_{2})|\geq 2^{-n}f(n)$. Further decomposing according to the last time $h\leq\log(n)^{1.5}$ at which $\widetilde{S}_{h,\widetilde{g}}(w)>\alpha\log(2)h$, we then express the pre-limit in the left-hand side of (3.8) as
$\sum_{j=1}^{Cf(n)^{2}}\sum_{h=0}^{\log(n)^{1.5}}\sum_{w\in\mathcal{N}_{n,j}(v)}|I_{n}(w)|\,\overline{M}_{n,\widetilde{a}}^{(1)}(w)\widetilde{E}_{n,\gamma,\widetilde{g}}\mathbbm{1}_{A_{h}^{\alpha}(w)}\mathbbm{1}_{B_{h,n}^{\alpha}(w)}\,.$
We now compute the second moment of the above. Below, we use $(\sum_{i=1}^{N}|x_{i}|)^{2}\leq N\sum_{i=1}^{N}|x_{i}|^{2}$ and the triangle inequality to pull the sums over $h$ and $j$ out of the square:
$\mathbb{E}\,\Bigg{|}\sum_{j=1}^{Cf(n)^{2}}\sum_{h=0}^{\log(n)^{1.5}}\sum_{w\in\mathcal{N}_{n,j}(v)}|I_{n}(w)|\,\overline{M}_{n,\widetilde{a}}^{(1)}(w)\widetilde{E}_{n,\gamma,\widetilde{g}}(w)\mathbbm{1}_{B_{\log(n)^{1.5},n}^{\alpha}(w)}\Bigg{|}^{2}$
$\qquad\leq Cf(n)^{2}\log(n)^{1.5}\sum_{j=1}^{Cf(n)^{2}}\sum_{h=0}^{\log(n)^{1.5}}\mathbb{E}\,\Bigg{|}\sum_{w\in\mathcal{N}_{n,j}(v)}|I_{n}(w)|\,\overline{M}_{n,\widetilde{a}}^{(1)}(w)\widetilde{E}_{n,\gamma,\widetilde{g}}(w)\mathbbm{1}_{A_{h}^{\alpha}(w)}\mathbbm{1}_{B_{\log(n)^{1.5},n}^{\alpha}(w)}\Bigg{|}^{2}$
@@
We begin by bounding $\mathsf{Diag}(h,j)$. Let $\mathcal{F}_{\ell}$ be the filtration including all random variables up to the $l$-th level in the tree. Conditioning on $\mathcal{F}_{h}$, then applying [Junnila, Lemma 10] and Lemma 2.2 yields that, for all $j\leq Cf(n)^{2}$, $h\leq\log(n)^{1.5}$, and $w\in\mathcal{N}_{n,j}(w)$, we have
@@
$\leq C\mathbb{E}\Big{[}\widetilde{E}_{h,\gamma,\widetilde{g}}(w)^{2}\Big{]}2^{(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})(n-h)}\leq C2^{\frac{\gamma^{2}}{2}n+\varepsilon(\alpha)n+\gamma^{2}\log(n)^{1.5}}\,,$
where $\varepsilon(\alpha)\searrow 0$ as $\alpha\searrow\gamma$, and we have taken $\alpha$ close to $\gamma$. This with Lemma 3.3 yields
$\max_{j\leq Cf(n)^{2},h\leq\log(n)^{1.5}}\mathsf{Diag}(h,j)$ $\leq C|\mathcal{N}_{n,j}(v)|2^{-2n}f(n)^{-2}\log(n)^{2}2^{\frac{\gamma^{2}}{2}n+\varepsilon(\alpha)n+\gamma^{2}\log(n)^{1.5}+\gamma^{2}\log(n)^{2}}$
@@
where in the second inequality we used $|\mathcal{N}_{n,j}(v)|\leq T_{n}$, followed by (2.1).
@@
$\mathcal{D}_{l,j}$ :=$:=\big{\{}(w_{1},w_{2})\in\mathcal{N}_{n,j}(v):|t(w_{1})-t(w_{2})|\in[2^{-l-1},2^{-l}]\big{\}}\,.$
Set $l_{0}:=l_{0}(n)=\min\{l\in\mathbb{N}:2^{-l}<2^{-n}f(n)\}$. By definition of $\mathcal{N}_{n,j}(v)$ we have
$\{w_{1}\neq w_{2}\in\mathcal{N}_{n,j}(v)\}=\bigsqcup_{l=3}^{l_{0}}\mathcal{D}_{l,j}.$
@@
~$\widetilde{\mathcal{B}}_{s,\gamma,\widetilde{g}}^{r,\alpha}(w):=\widetilde{E}_{s,\gamma,\widetilde{g}}(w)\mathbbm{1}_{A_{r}^{\alpha}(w)\cap B_{r,s}^{\alpha}(w)}\,.$
@@
We first consider summands with $l\geq h$ in (3.16) (note for $l>\log(n)^{1.5}$, this is automatic). For any $(w_{1},w_{2})\in\mathcal{D}_{\ell,j}$, we find
@@
$=\mathbb{E}\Big{[}\widetilde{\mathcal{B}}_{\ell,\gamma,\widetilde{g}}^{h,\alpha}(w_{1})\widetilde{\mathcal{B}}_{\ell,\gamma,\widetilde{g}}^{h,\alpha}(w_{2})\,\mathbb{E}\Big{[}\widetilde{E}_{(l,n],\gamma,\widetilde{g}}(w_{1})\widetilde{E}_{(l,n],\gamma,\widetilde{g}}(w_{2})\mathbbm{1}_{B_{l+1,n}^{\alpha}(w_{1})\cap B_{l+1,n}^{\alpha}(w_{2})}|\mathcal{F}_{l}\Big{]}\Big{]}$
@@
where the last step uses [Junnila, Lemma 9] and $|t_{n}(w_{1})-t_{n}(w_{2})|\leq 2^{-h}$. Conditioning on $\mathcal{F}_{h}$ then applying [Junnila, Lemma 10] gives
$\mathbb{E}\Big{[}\widetilde{\mathcal{B}}_{\ell,\gamma,\widetilde{g}}^{h,\alpha}(w_{1})\widetilde{\mathcal{B}}_{\ell,\gamma,\widetilde{g}}^{h,\alpha}(w_{2})\Big{]}$ $\leq\mathbb{E}\Big{[}\widetilde{E}_{h,\gamma,\widetilde{g}}(w_{1})\widetilde{E}_{h,\gamma,\widetilde{g}}(w_{2})\Big{]}2^{(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})(l-h)}\,.$ (3.18)
Lemma 2.2 bounds the right-hand side of the above display by
$C2^{\gamma^{2}h}2^{(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})(l-h)}\leq C2^{\gamma^{2}\log(n)^{1.5}+(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})l}\,.$ (3.19)
Next, we consider summands in (3.16) with $l\leq h-1$. We proceed as in Sections 3.2.2, 3.18 and 3.19, this time conditioning on $\mathcal{F}_{h+1}$ and using $|t_{n}(w_{1})-t_{n}(w_{2})|\geq 2^{-l-1}\geq 2^{-h-1}$ in [Junnila, Lemma 9]:
@@
Combining (3.2.2)–(3.20) yields that, for all $h\leq\log(n)^{1.5}$, $j\leq Cf(n)^{2}$, $(w_{1},w_{2})\in\mathcal{D}_{l,j}$, and $\ell\in[3,\ell_{0}]$,
$\mathbb{E}\Big{[}\widetilde{\mathcal{B}}_{n,\gamma,\widetilde{g}}^{h,\alpha}(w_{1})\widetilde{\mathcal{B}}_{n,\gamma,\widetilde{g}}^{h,\alpha}(w_{2})\Big{]}\leq C2^{\gamma^{2}\log(n)^{1.5}+(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})l}\,.$ (3.21)
Substituting (3.21) and (2.1) into (3.16) yields, for $j\leq Cf(n)^{2},h\leq\log(n)^{1.5}$ arbitrary
$\mathsf{Cross}(h,j)\leq C\sum_{l=3}^{l_{0}}\sum_{(w_{1},w_{2})\in\mathcal{D}_{l,j}}2^{-2n}f(n)^{-2}\mathbb{E}\Big{[}\overline{M}_{n,\widetilde{a}}^{(1)}(w_{1})\overline{M}_{n,\widetilde{a}}^{(1)}(w_{2})\Big{]}2^{\gamma^{2}\log(n)^{1.5}}2^{(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})l}\,.$
Observe for $l\leq\log(n)^{2}$ and $(w_{1},w_{2})\in\mathcal{D}_{l,j}$ we have
@@
for $n$ sufficiently big, so that Lemma 3.5 yields
[…]
Furthermore, by Lemma B.1, we have $|\mathcal{D}_{l,j}|\leq T_{n}^{2}2^{-l}=2^{2n}f(n)^{2}2^{-l}$, which yields
@@
In light of (3.2.2), choose also $\alpha$ close enough to $\gamma$ so that $\frac{\gamma^{2}}{2}+\varepsilon(\alpha)-1<0$, which is possible since $\gamma<\sqrt{2}$. Then substituting (3.2.2), (3.22), and Lemma 3.3 into (3.2.2) yields
@@
We use the last display to conclude (3.8) by Markov’s inequality and Borel-Cantelli. ∎
3.2.3. Proofs of Lemmas 3.3, 3.4 and 3.5
Towards Eq. 3.2, it remains to prove Lemmas 3.3, 3.4 and 3.5.
@@
It is enough to show the bound for $\mathbb{E}[M_{n,\widetilde{a}}^{(1)}(w)^{2}]$, since the first moment is bounded by 1 by comparison to $M_{n,\widetilde{a}}(w)$. We plan to decompose $M_{n,\widetilde{a}}^{(1)}(w)$ according to the last generation $h$ (counting forward from $n$) at which the field at $t$ was more than $\alpha$ thick. Roughly speaking, the only contribution will come from the value of the field at generation $h$. For this purpose, we set $h_{n}:=n+h+1$ and
$A_{h;n}^{\alpha}(t)$ :=$:=\big{\{}\widetilde{S}_{(n,n+h],\widetilde{a}}(t)\geq\alpha\log(2)h\big{\}}\,.$ (3.23)
@@
$M_{n,\widetilde{a}}^{(1)}(w)$ $=\lim\limits_{r\to\infty}|I_{n}(w)|^{-1}\sum_{h=0}^{\log(n)^{2}}\int_{I_{n}(w)}\mathbbm{1}_{A_{h;n}^{\alpha}(t)\cap B_{h,r;n}^{\alpha}(t)}e^{\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(t)}\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(t)^{-1}\;\mathrm{d}t$
$=\lim\limits_{r\to\infty}|I_{n}(w)|^{-1}\sum_{h=0}^{\log(n)^{2}}\sum_{u\in\mathcal{N}_{h_{n}}(w)}\int_{I_{h_{n}}(u)}\mathbbm{1}_{A_{h;n}^{\alpha}(t)\cap B_{h,r;n}^{\alpha}(t)}e^{\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(t)}\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(t)^{-1}\;\mathrm{d}t,$
where in the second step we have decomposed $I_{n}(w)$ into its subintervals in generation $h_{n}$. Applying Fatou and Fubini now yields
$\mathbb{E}\Big{[}M_{n,\widetilde{a}}^{(1)}(w)^{2}\Big{]}\leq\liminf\limits_{r\to\infty}|I_{n}(w)|^{-2}\log(n)^{2}\\ \times\max_{h\leq\log(n)^{2},u\in\mathcal{N}_{h_{n}}(w)}|\mathcal{N}_{h_{n}}(w)|^{2}\,\mathbb{E}\bigg{|}\int_{I_{h_{n}}(u)}\mathbbm{1}_{A_{h;n}^{\alpha}(t)\cap B_{h,r;n}^{\alpha}(t)}e^{\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(t)}\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(t)^{-1}\mathrm{d}t\bigg{|}^{2}\,.$ (3.24)
Let $\mathfrak{T}(w)$ denote the subtree of $\mathfrak{T}$ below $w$. In our embedding of $\mathfrak{T}$ into $[0,1]$, the subtree $\mathfrak{T}(w)$ embeds into the interval $I_{n}(w)$ of length (roughly (2.1)) $2^{-n}f(n)^{-1}$. We now embed $\mathfrak{T}(w)$ into $[0,1]$ by scaling space; this way $(\widetilde{S}_{(n,n+r]}(t))_{r\in\mathbb{N},t\in[0,1]}$ is $\log$-correlated in the sense of Remark 2.1, and in particular the results of [Junnila] apply. Letting $\lambda_{w}$ denote the rescaled Lebesgue measure, we have $\lambda_{w}(I_{h_{n}})\leq 2^{-h-1}$. We now apply [Junnila, Proposition 11] to the second moment on the right-hand side of (3.24) to see that, for all $h\leq\log(n)^{2}$and $u\in\mathcal{N}_{h_{n}}(w)$, we have
$\mathbb{E}\bigg{|}\int_{I_{h_{n}}(u)}\mathbbm{1}_{A_{h;n}^{\alpha}(t)\cap B_{h,r;n}^{\alpha}(t)}e^{\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(t)}\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(t)^{-1}\;\mathrm{d}t\bigg{|}^{2}\\ \leq 2^{-2h}|I_{n}(w)|^{2}\mathbb{E}\bigg{[}\sup_{t\in I_{h_{n}}(u)}\Big{(}e^{\gamma\widetilde{S}_{(n,n+h],\widetilde{a}}(t)}\widetilde{Z}_{(n,n+h],\gamma,\widetilde{a}}(t)^{-1}\Big{)}^{2}\bigg{]}\,.$ (3.25)
By definition of the model, $\widetilde{S}_{(n,n+h],\widetilde{a}}(t)=\widetilde{S}_{(n,n+h],\widetilde{a}}(u)$ is constant for all $t\in I_{n+h+1}(u)$, which together with Lemma 2.2 yields
$2^{-2h}|I_{n}(w)|^{2}\mathbb{E}\bigg{[}\sup_{t\in I_{n+h+1}(u)}\left(e^{\gamma\widetilde{S}_{(n,n+h],\widetilde{a}}(t)}\widetilde{Z}_{(n,n+h],\gamma,\widetilde{a}}(t)^{-1}\right)^{2}\bigg{]}\leq C2^{-2h+\gamma^{2}h}|I_{n}(w)|^{2}\,.$ (3.26)
Plugging (3.25) and (3.26) into (3.24) yields
@@
The lemma then follows from the bound on $|\mathcal{N}_{h_{n}}(w)|$ provided by Eq. B.1. ∎

### v2

@@ 3.2.1. Reduction, and overview of the proof
in such a way that any $w_{1}\neq w_{2}$ in $\mathcal{N}_{n,j}(v)$ satisfies $1/8\geq|t_{n}(w_{1})-t_{n}(w_{2})|\geq 2^{-n}f(n)$. Further decomposing according to the last level $h\leq\log(n)^{1.5}$ at which $\widetilde{S}_{h,\widetilde{g}}(w)>\alpha\log(2)h$, we then express the pre-limit in the left-hand side of Eq.˜3.8 as
$\sum_{j=1}^{Cf(n)^{2}}\sum_{h=0}^{\log(n)^{1.5}}\sum_{w\in\mathcal{N}_{n,j}(v)}|I_{n}(w)|\,\overline{M}_{n,\widetilde{a}}^{(1)}(w)\widetilde{E}_{n,\gamma,\widetilde{g}}(w)\mathbbm{1}_{A_{h}^{\alpha}(w)}\mathbbm{1}_{B_{h,n}^{\alpha}(w)}\,.$
We now compute the second moment of the above. Below, we use the fact $(\sum_{i=1}^{N}|x_{i}|)^{2}\leq N\sum_{i=1}^{N}|x_{i}|^{2}$ and the triangle inequality to pull the sums over $h$ and $j$ out of the square:
$\mathbb{E}\,\Bigg|\sum_{j=1}^{Cf(n)^{2}}\sum_{h=0}^{\log(n)^{1.5}}\sum_{w\in\mathcal{N}_{n,j}(v)}|I_{n}(w)|\,\overline{M}_{n,\widetilde{a}}^{(1)}(w)\widetilde{E}_{n,\gamma,\widetilde{g}}(w)\mathbbm{1}_{A_{h}^{\alpha}(w)}\mathbbm{1}_{B_{h,n}^{\alpha}(w)}\Bigg|^{2}$
$\qquad\leq Cf(n)^{2}\log(n)^{1.5}$
$\quad\qquad\times\sum_{j=1}^{Cf(n)^{2}}\sum_{h=0}^{\log(n)^{1.5}}\mathbb{E}\,\Bigg|\sum_{w\in\mathcal{N}_{n,j}(v)}|I_{n}(w)|\,\overline{M}_{n,\widetilde{a}}^{(1)}(w)\widetilde{E}_{n,\gamma,\widetilde{g}}(w)\mathbbm{1}_{A_{h}^{\alpha}(w)}\mathbbm{1}_{B_{h,n}^{\alpha}(w)}\Bigg|^{2}$
@@
We begin by bounding $\mathsf{Diag}(h,j)$. Let $\mathcal{F}_{\ell}$ be the filtration generated by all random variables up to the $l^{\mathrm{th}}$ level in the tree. Conditioning on $\mathcal{F}_{h}$, then applying [25, Lemma 10] and Lemma˜2.2 yields that, for all $j\leq Cf(n)^{2}$, $h\leq\log(n)^{1.5}$, and $w\in\mathcal{N}_{n,j}(w)$, we have
@@
$\leq C\mathbb{E}\Big[\widetilde{E}_{h,\gamma,\widetilde{g}}(w)^{2}\Big]2^{(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})(n-h)}$
$\leq C2^{\frac{\gamma^{2}}{2}n+\varepsilon(\alpha)n+\gamma^{2}\log(n)^{1.5}}\,$,
where $\varepsilon(\alpha)\searrow 0$ as $\alpha\searrow\gamma$, and we have taken $\alpha$ close to $\gamma$. This with Lemma˜3.3 yields
$\max_{j\leq Cf(n)^{2},h\leq\log(n)^{1.5}}\mathsf{Diag}(h,j)$ $\leq C|\mathcal{N}_{n,j}(v)|2^{-2n}f(n)^{-2}\log(n)^{2}2^{\frac{\gamma^{2}}{2}n+\varepsilon(\alpha)n+2\gamma^{2}\log(n)^{2}}$
@@
where in the second inequality we used $|\mathcal{N}_{n,j}(v)|\leq T_{n}$, followed by Eq.˜2.1.
@@
$\mathcal{D}_{\ell,j}$ $:=\big\{(w_{1},w_{2})\in\mathcal{N}_{n,j}(v):|t(w_{1})-t(w_{2})|\in[2^{-\ell-1},2^{-\ell}]\big\}\,.$
Set $\ell_{0}:=\ell_{0}(n)=\min\{\ell\in\mathbb{N}:2^{-\ell}<2^{-n}f(n)\}$. By definition of $\mathcal{N}_{n,j}(v)$ we have
$\{w_{1}\neq w_{2}\in\mathcal{N}_{n,j}(v)\}=\bigsqcup_{\ell=3}^{\ell_{0}}\mathcal{D}_{l,j}.$
@@
$\widetilde{\mathcal{B}}_{s,\gamma,\widetilde{g}}^{r,\alpha}(w):=\widetilde{E}_{s,\gamma,\widetilde{g}}(w)\mathbbm{1}_{A_{r}^{\alpha}(w)\cap B_{r,s}^{\alpha}(w)}\,.$
@@
We first consider summands with $\ell\geq h$ in Eq.˜3.16 (note for $\ell>\log(n)^{1.5}$, this is automatic). For any $(w_{1},w_{2})\in\mathcal{D}_{\ell,j}$, we find
@@
$=\mathbb{E}\Big[\widetilde{\mathcal{B}}_{\ell,\gamma,\widetilde{g}}^{h,\alpha}(w_{1})\widetilde{\mathcal{B}}_{\ell,\gamma,\widetilde{g}}^{h,\alpha}(w_{2})\,\mathbb{E}\Big[\widetilde{E}_{(\ell,n],\gamma,\widetilde{g}}(w_{1})\widetilde{E}_{(\ell,n],\gamma,\widetilde{g}}(w_{2})\mathbbm{1}_{B_{\ell+1,n}^{\alpha}(w_{1})\cap B_{l+1,n}^{\alpha}(w_{2})}|\mathcal{F}_{\ell}\Big]\Big]$
@@
where the last step uses [25, Lemma 9] and $|t_{n}(w_{1})-t_{n}(w_{2})|\leq 2^{-h}$. Conditioning on $\mathcal{F}_{h}$ then applying [25, Lemma 10] gives
$\mathbb{E}\Big[\widetilde{\mathcal{B}}_{\ell,\gamma,\widetilde{g}}^{h,\alpha}(w_{1})\widetilde{\mathcal{B}}_{\ell,\gamma,\widetilde{g}}^{h,\alpha}(w_{2})\Big]$ $\leq\mathbb{E}\Big[\widetilde{E}_{h,\gamma,\widetilde{g}}(w_{1})\widetilde{E}_{h,\gamma,\widetilde{g}}(w_{2})\Big]2^{(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})(\ell-h)}\,.$ (3.18)
Lemma˜2.2 bounds the right-hand side of the above display by
$C2^{\gamma^{2}h}2^{(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})(\ell-h)}\leq C2^{\gamma^{2}\log(n)^{1.5}+(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})\ell}\,.$ (3.19)
Next, we consider summands in Eq.˜3.16 with $\ell\leq h-1$. We proceed as in Sections˜3.2.2, 3.18 and 3.19, this time conditioning on $\mathcal{F}_{h+1}$ and using $|t_{n}(w_{1})-t_{n}(w_{2})|\geq 2^{-\ell-1}\geq 2^{-h-2}$ in [25, Lemma 9]:
@@
Combining Section˜3.2.2Eq.˜3.20 yields that, for all $h\leq\log(n)^{1.5}$, $j\leq Cf(n)^{2}$, $(w_{1},w_{2})\in\mathcal{D}_{\ell,j}$, and $\ell\in[3,\ell_{0}]$,
$\mathbb{E}\Big[\widetilde{\mathcal{B}}_{n,\gamma,\widetilde{g}}^{h,\alpha}(w_{1})\widetilde{\mathcal{B}}_{n,\gamma,\widetilde{g}}^{h,\alpha}(w_{2})\Big]\leq C2^{\gamma^{2}\log(n)^{1.5}+(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})\ell}\,.$ (3.21)
Substituting Eq.˜3.21 and Eq.˜2.1 into Eq.˜3.16 yields, for $j\leq Cf(n)^{2},h\leq\log(n)^{1.5}$ arbitrary
$\mathsf{Cross}(h,j)\\ \leq C\sum_{\ell=3}^{\ell_{0}}\sum_{(w_{1},w_{2})\in\mathcal{D}_{\ell,j}}2^{-2n}f(n)^{-2}\mathbb{E}\Big[\overline{M}_{n,\widetilde{a}}^{(1)}(w_{1})\overline{M}_{n,\widetilde{a}}^{(1)}(w_{2})\Big]2^{\gamma^{2}\log(n)^{1.5}}2^{(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})\ell}\,.$
Observe for $\ell\leq\log(n)^{2}$ and $(w_{1},w_{2})\in\mathcal{D}_{\ell,j}$ we have
@@
for $n$ sufficiently big, so that Lemma˜3.5 yields
$\mathsf{Cross}(h,j)$ $\leq C\sum_{\ell=3}^{\log(n)^{2}}|\mathcal{D}_{\ell,j}|\,2^{-2n}f(n)^{-2}2^{-\frac{(\alpha-\gamma^{2})}{2}\log(n)^{2}+\gamma^{2}\log(n)^{1.5}+(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})\ell}$
$\qquad+C\sum_{\ell=\log(n)^{2}+1}^{\ell_{0}}|\mathcal{D}_{\ell,j}|\,2^{-2n}f(n)^{-2}n^{-2}2^{\gamma^{2}\log(n)^{1.5}+(\gamma^{2}-\frac{(2\gamma-\alpha)^{2}}{2})\ell}\,.$
Furthermore, by Lemma B.1, we have $|\mathcal{D}_{\ell,j}|\leq T_{n}^{2}2^{-\ell}=2^{2n}f(n)^{2}2^{-\ell}$, which yields
@@
In light of Section˜3.2.2, choose also $\alpha$ close enough to $\gamma$ so that $\frac{\gamma^{2}}{2}+\varepsilon(\alpha)-1<0$, which is possible since $\gamma<\sqrt{2}$. Then substituting Section˜3.2.2, Eq.˜3.22, and Lemma 3.3 into Section˜3.2.2 yields
@@
We use the last display to conclude Eq.˜3.8 by Markov’s inequality and Borel-Cantelli. ∎
3.2.3. Proofs of Lemmas˜3.3, 3.4 and 3.5
Towards Eq.˜3.2, it remains to prove Lemmas˜3.3, 3.4 and 3.5.
@@
It is enough to show the bound for $\mathbb{E}[M_{n,\widetilde{a}}^{(1)}(w)^{2}]$, since the first moment is bounded by 1 by comparison to $M_{n,\widetilde{a}}(w)$. We plan to decompose $M_{n,\widetilde{a}}^{(1)}(w)$ according to the last generation $h$ (counting forward from $n$) at which the field at $t$ was more than $\alpha$ thick. Roughly speaking, the only contribution comes from the value of the field at generation $h$. For this purpose, we set $h_{n}:=n+h+1$ and
$A_{h;n}^{\alpha}(t)$ $=\big\{\widetilde{S}_{(n,n+h],\widetilde{a}}(t)>\alpha\log(2)h\big\}\,.$ (3.23)
@@
$M_{n,\widetilde{a}}^{(1)}(w)$ $=\lim\limits_{r\to\infty}|I_{n}(w)|^{-1}\sum_{h=0}^{\log(n)^{2}}\int_{I_{n}(w)}\mathbbm{1}_{A_{h;n}^{\alpha}(t)\cap B_{h,r;n}^{\alpha}(t)}\widetilde{E}_{(n,n+r],\gamma,\widetilde{a}}(v_{n+r}(t))\;\mathrm{d}t$
$=\lim\limits_{r\to\infty}|I_{n}(w)|^{-1}\sum_{h=0}^{\log(n)^{2}}\sum_{u\in\mathcal{N}_{h_{n}}(w)}\int_{I_{h_{n}}(u)}\mathbbm{1}_{A_{h;n}^{\alpha}(t)\cap B_{h,r;n}^{\alpha}(t)}\widetilde{E}_{(n,n+r],\gamma,\widetilde{a}}(v_{n+r}(t))\;\mathrm{d}t,$
where in the second step we have decomposed $I_{n}(w)$ into its subintervals in generation $h_{n}$. Applying Fatou’s lemma and Fubini’s theorem now yields
$\mathbb{E}\Big[M_{n,\widetilde{a}}^{(1)}(w)^{2}\Big]\leq\liminf\limits_{r\to\infty}|I_{n}(w)|^{-2}\log(n)^{2}\\ \times\max_{h\leq\log(n)^{2},u\in\mathcal{N}_{h_{n}}(w)}|\mathcal{N}_{h_{n}}(w)|^{2}\,\mathbb{E}\bigg|\int_{I_{h_{n}}(u)}\mathbbm{1}_{A_{h;n}^{\alpha}(t)\cap B_{h,r;n}^{\alpha}(t)}\widetilde{E}_{(n,n+r],\gamma,\widetilde{a}}(v_{n+r}(t))\mathrm{d}t\bigg|^{2}\,.$ (3.24)
Let $\mathfrak{T}(w)$ denote the subtree of $\mathfrak{T}$ below $w$. In our embedding of $\mathfrak{T}$ into $[0,1]$, the subtree $\mathfrak{T}(w)$ embeds into the interval $I_{n}(w)$ of length (roughly Eq.˜2.1) $2^{-n}f(n)^{-1}$. We now embed $\mathfrak{T}(w)$ into $[0,1]$ by scaling space; this way $(\widetilde{S}_{(n,n+r]}(t))_{r\in\mathbb{N},t\in[0,1]}$ is $\log$-correlated in the sense of Remark˜2.1, and in particular the results of [25] apply. Letting $\lambda_{w}$ denote the rescaled Lebesgue measure, we have $\lambda_{w}(I_{h_{n}})\leq 2^{-h-1}$. We now apply [25, Proposition 11] to the second moment on the right-hand of Eq.˜3.24 to see that, for all $h\leq\log(n)^{2}$and $u\in\mathcal{N}_{h_{n}}(w)$,
$\mathbb{E}\bigg|\int_{I_{h_{n}}(u)}\mathbbm{1}_{A_{h;n}^{\alpha}(t)\cap B_{h,r;n}^{\alpha}(t)}e^{\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(t)}\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(t)^{-1}\;\mathrm{d}t\bigg|^{2}\\ \leq 2^{-2h}|I_{n}(w)|^{2}\mathbb{E}\bigg[\sup_{t\in I_{h_{n}}(u)}e^{2\gamma\widetilde{S}_{(n,n+h],\widetilde{a}}(t)}\widetilde{Z}_{(n,n+h],\gamma,\widetilde{a}}(t)^{-2}\bigg]\,.$ (3.25)
By definition of the model, $\widetilde{S}_{(n,n+h],\widetilde{a}}(t)=\widetilde{S}_{(n,n+h],\widetilde{a}}(u)$ is constant for all $t\in I_{h_{n}}(u)$, which together with Lemma˜2.2 yields
$2^{-2h}|I_{n}(w)|^{2}\mathbb{E}\bigg[\sup_{t\in I_{n+h+1}(u)}e^{2\gamma\widetilde{S}_{(n,n+h],\widetilde{a}}(t)}\widetilde{Z}_{(n,n+h],\gamma,\widetilde{a}}(t)^{-2}\bigg]\leq C2^{-2h+\gamma^{2}h}|I_{n}(w)|^{2}\,.$ (3.26)
Plugging Eq.˜3.25 and Eq.˜3.26 into Eq.˜3.24 yields
@@
The lemma then follows from the bound on $|\mathcal{N}_{h_{n}}(w)|$ provided by Eq.˜B.1. ∎